### v3

@@ 3.2 Weighted Voters and Maynard-Cross Learning
We now consider a large population of $N\gg 1$ honey bees reaching a consensus on which nesting site to select by applying $R_{\text{wvoter}}$. In this Section, we show that, although $R_{\text{wvoter}}$ is a blind imitation protocol where individual bees simply imitate the first type they encounter, a swarm of bees following $R_{\text{wvoter}}$ aggregates to an RL agent at the macro-organism level. In the $R_{\text{wvoter}}$ model, individual bees are blind imitators, but active promoters. Each bee $i$ of type $T_{i}=k$ and payoff sample $r_{k}\sim r(k)$ has a tangible stochastic influence on the inflow of other bees that it rallies to its own type $k$. We represent this influence by the corresponding change in the $k^{\text{th}}$ component of the population vector, which we denote $P^{(i)}(k\leftarrow\cdot)$:
$P^{(i)}(k\leftarrow\cdot)=\frac{r_{k}}{\sum_{j}r^{(j)}},$ (12)
where $r_{k}$ represents $i$’s broadcasting frequency, and $\sum_{j}r^{(j)}$ represents the total broadcasting frequency of the entire swarm. Thus, in a well-mixed population, the outflow from any type $a\neq k$ to the type $k$ attributable to $i$ is:
$P^{(i)}(k\leftarrow a)$ $=\pi_{a}\frac{r_{k}}{\sum_{j}r^{(j)}}$
$=\frac{1}{N}\pi_{a}\frac{r_{k}}{\frac{1}{N}\sum_{j}r^{(j)}}$
@@
Here, the meaning of $\alpha$ is of peculiar interest. In Eq. 15, $\alpha$ looks very much like a learning rate, i.e., a hyperparameter that RL practitioners typically set to a small value in order to downsize the amplitude of individual policy updates. When dealing with on-policy RL algorithms, computer scientists often evaluate a policy in several parallel copies of the same environment, average the corresponding updates into a single policy update, and downsize this update via a small learning rate. The common rule of thumb is that the learning rate can be larger when there are enough parallel environments. Instead, in Definition 3.2, $\alpha$ has a clear population-based meaning: it is the inverse of the number of bees acting in parallel in the swarm, as we are describing the influence of one single bee on the population. Crucially, this influence is that of an RL agent, but the individual bee itself is in no way an RL agent: it only blindly imitates its peers! Instead, everything happens as if nature had evolved individual bees’ behavior such that it aggregates to an RL agent at the macro-organism level: the “hive mind”, whose rewards are the quality estimates of nest site samples. Summing the influence of all individual bees yields:
@@
Proof 3.3.
Direct consequence of Eqs. 16 and 17.
4 Discussion
@@ 4.1 A single RL agent in many parallel environments
Propositions 1 and 2 draw a formal description of collective intelligence. In particular, researchers in Swarm Intelligence [swarmintelligence99] have already informally witnessed this type of intelligence [franks2002information], often referred to as “emergent” to describe situations where simple, seemingly blind individual behaviors aggregate to a complex and coherent behavior of the group. 2 shows that, when the individual bees’ behavior is modeled as $R_{\text{wvoter}}$, this “emergent” collective intelligence is MCL, a multi-armed bandit Reinforcement Learning algorithm. But the implications of 2 go even further: not only a swarm of bees is a single MCL agent, it is one that is quite efficient at what it does, because each bee is a parallel action sample of its policy. In other words, each bee $i$ can be seen as an action sample $a^{(i)}\sim\pi$ tested against a parallel copy $E^{(i)}$ of the environment $E$, as illustrated in Fig. 1.
Figure 2: $\alpha$-MCL for varying learning rates
@@
Fig. 2 shows that this converges much faster than if individual bees were themselves to learn via $\alpha$-MCL instead of just following $R_{\text{wvoter}}$. If the MCL algorithm were of any use outside of biology [^2] Most likely not: more advanced algorithms exist to solve multi-armed bandits (UCB…)., say if a computer scientist wanted to train a robot via iterative $\alpha$-MCL, they would need a small $\alpha$ for the algorithm to converge to the optimal option. Or, if feasible, they could afford using a larger $\alpha$ by evaluating the policy in several parallel simulations (Sections 7.5 and 7.6). As we saw, a swarm of honey bees naturally does this in as many environments as there are scouts in the swarm. Finally, 1 implies that a similar discussion holds for the “imitation of success” revision protocol, which encompasses certain human behaviors, and even natural selection itself.
4.2 Importance of the swarm size
Figure 3: $R_{\text{wvoter}}$ swarms of varying sizes $N$
@@ 4.2 Importance of the swarm size
In Section 4.1, we have discussed how being made of a number of individuals enables the “hive mind” to evaluate actions sampled from $\pi$ in a parallel fashion. Furthermore, our analysis assumed that this number was large (which enabled us to compute flows in Section 3). Yet, in practice there is a tradeoff here. Actual honey bee swarms are known to use only a small proportion of their bees as scouts, while the large majority of the swarm remains quiescent during nest-site selection [seeley1978nest, beekman2006does], which is surmised to be in the interest of energy-efficiency. In a typical swarm of 10,000 bees, roughly 200 to 500 scouts are active during nest-site selection. To analyze how the number of scouts $N$ affects the task-performance of the MCL macro-agent, we conducted simplified simulations where swarms of varying sizes had to chose the best amongst 10 options. Option qualities were spread between 0 and 1, while the noise in bees’ estimates was modeled by a uniform perturbation of amplitude $0.2$. Fig. 3 reports the percentage of populations that converged to the optimal choice for different values of $N$, across 1000 seeds (each individual seed converged to a homogeneous choice). In these simulations, $N=500$ is enough to closely follow the MRD and converge to the optimal option, whereas a number of scout bees that is too small often yields convergence to a sub-optimal nest site option. It is however important to stress that our choice of simulation parameters is largely arbitrary: Fig. 3 is qualitative rather than quantitative.
4.3 Blind imitation over imitation of success
@@
The $R_{wvoter}$ revision protocol adopted by honey bees has surprising advantages over imitation of success. In Section 7, we show that $R_{wvoter}$ (MCL) generally converges faster than $R_{success}$ (CL), except for small population sizes. This is in addition to the simplicity of $R_{wvoter}$, which require bees to just blindly mimic their peers.
@@ 4.4 Broader significance
  • Biology. 2 shows that seemingly blind, imitation-based bee interactions aggregate to a form of collective intelligence, which is that of a “hive mind” learning through reinforcement. This is interesting from a neuroscience perspective, since both imitation and reinforcement have long been known to govern learning at the individual level [Zentall2006, neftci2019reinforcement]. What our paper shows about honey bees is that their puzzling imitative behavior can in fact be understood as reinforcement from the swarm-agent perspective, where this behavior equates to massively parallel MCL. Why this imitative behavior has evolved in nature thus seems clearer: swarms of honey bees are an example of reinforcement strategy having evolved at the group level. This analysis agrees with the commonly accepted theory of group/kin selection to explain the behavior of honey bees [robinson89].
  • Economy and Social Dynamics. Evolutionary Game Theory goes beyond modeling biology: it extends to many types of population dynamics, most notably in economics [borgers1997learning, apesteguia2007imitation]. When thinking of $R_{\text{success}}$ as a model of certain decision-making strategies (e.g., trust in online reviews), 1 yields that imitating others’ success makes us part of a macroscopic RL process that hugely benefits the whole group (see Section 4.1). This opens an avenue for modeling the group-level impacts of economic actors’ imitative behaviors as forms of collective reinforcement learning.
  • Algorithmic Swarm Intelligence. The field of Swarm Intelligence takes inspiration from complex emergent behaviors arising from a collective of natural entities following simple, local, and decentralized rules [swarmintelligence99], to engineer algorithms that leverage collective principles such as coordination, cooperation, and communication to tackle various problems (including Collective-Decsision Making) across multiple domains such as swarm robotics [swarmdorigo2021, heikobook18] and optimization [dorigoACO06]. Propositions 1 and 2 provide theoretical grounding for the phenomenon commonly referred to as “emergence” in the field of SI. This bridge also opens up the possibility of porting ideas from Reinforcement Learning to derive Swarm Intelligence local rules.
  • Algorithmic Reinforcement Learning. Similar to how we abstracted a population of $R_{wvoter}$ individuals as a single MCL agent, other revision models (majority rule [valentini2016collective], cross-inhibition [Reina_2017]…) can be used to abstract dynamic groups of entities as simple, single RL agents. In massively multi-agent real-world settings (e.g., finance), this could help model the non-stationarity of environment dynamics.
@@ 4.5 Conclusion
In 1 and 2, we formalized equivalences between imitation-based revision models in large populations, and reinforcement learning when the said populations are seen as single coherent agents. Because the models considered in this paper are rooted in biology and social dynamics, our analysis has deep real world implications. In particular, 2 shows that the “hive mind” that biologists observed in swarms of bees can indeed be seen as a single coherent agent learning via reinforcement. In the simplified model of weighted voters, this agent learns via Maynard-Cross Learning.

### v4

@@ 3.2 Weighted Voters and Maynard-Cross Learning
We now consider a large population of $N\gg 1$ honey bees with average local neighborhood size $M\gg 1$, seeking agreement on which nesting site to select by applying $R_{\text{wvoter}}$. We show that, although $R_{\text{wvoter}}$ is a blind imitation protocol where individual bees simply imitate the first type they encounter, a swarm of bees following $R_{\text{wvoter}}$ aggregates to an RL agent at the macro-organism level. In the $R_{\text{wvoter}}$ model, individual bees are blind imitators, but active promoters. Each bee $i$ of type $T_{i}=k$ and payoff sample $r^{(i)}_{k}\sim r(k)$ has a tangible stochastic influence on the local expected inflow of other bees that it rallies to its own type $k$ within its local neighborhood $\mathcal{N}_{i}$:
$P^{(i)}_{\text{neighborhood}}(k\leftarrow\cdot)=\frac{r^{(i)}_{k}}{\sum\limits_{j\in\mathcal{N}_{i}}r^{(j)}},$ (8)
where $r^{(i)}_{k}$ represents $i$’s broadcasting frequency, and $\sum_{j\in\mathcal{N}_{i}}r^{(j)}$ represents the total broadcasting frequency of $i$’s local neighborhood. The expected outflow attributable to $i$ on the entire swarm from any type $a\neq k$ to the type $k$ is thus:
$P^{(i)}(k\leftarrow a)=\frac{N^{(i)}_{a}}{N}\frac{r^{(i)}_{k}}{\sum\limits_{j\in\mathcal{N}_{i}}r^{(j)}},$ (9)
where $N^{(i)}_{a}$ is the number of type-$a$ bees within $i$’s neighborhood. For simplicity, we assume that $M$ (and thus also $N^{(i)}$) is large. Assuming that all individuals are well mixed in the population [31], it follows that $\frac{N^{(i)}_{a}}{M}=\pi_{a}$:
$P^{(i)}(k\leftarrow a)$ $=\frac{M\pi_{a}}{N}\frac{r^{(i)}_{k}}{\sum\limits_{j\in\mathcal{N}_{i}}r^{(j)}}$
$=\frac{\pi_{a}}{N}\frac{r_{k}}{\frac{1}{M}\sum\limits_{j\in\mathcal{N}_{i}}r^{(j)}}$
@@
Here, the meaning of $\alpha$ is of peculiar interest. In Eq. 12, $\alpha$ looks very much like a learning rate, i.e., a hyperparameter that RL practitioners typically set to a small value in order to downsize the amplitude of individual policy updates. When dealing with on-policy RL algorithms, computer scientists often evaluate a policy in several parallel copies of the same environment [34], average the corresponding updates into a single policy update, and downsize this update via a small learning rate. The common rule of thumb is that the learning rate can be larger when there are enough parallel environments. But in Definition 4, $\alpha$ instead has a clear population-based meaning: it is the inverse of the number of bees acting in parallel in the swarm, as we are describing the influence of one single bee on the population. Crucially, this influence is that of an RL agent, but the individual bee itself is in no way an RL agent: it only blindly imitates its peers! Instead, everything happens as if nature had evolved individual bees’ behavior such that it aggregates to an RL agent at the macro-organism level: the “hive mind”, whose rewards are the quality estimates of nest site samples. Summing the influence of all individual bees yields:
@@
Proof.
Direct consequence of Eqs. 13 and 14. ∎
4 Final remarks and simulations
(a) $\alpha$-MCL for varying learning rates
(b) Varying swarm sizes NN
(c) Varying neighborhood sizes for NN=500
(d) Variants of $R_{\text{success}}$ & $R_{\text{wvoter}}$ in NN=1000
Figure 2: Simulations: (a) Swarms of bees reach consensus more rapidly when following $R_{\text{wvoter}}$ collectively than when individual bees learn via $\alpha$-MCL. (b,c) Varying swarm and neighborhood sizes show that the theoretical predictions hold under practical constraints. (d) Simple variants of $R_{\text{success}}$ and $R_{\text{wvoter}}$ can surpass $R_{\text{wvoter}}$ in this environment, raising the question of why evolution favored $R_{\text{wvoter}}$ over alternative decision-making strategies.
@@ 4.1 A single RL agent in many parallel environments
Researchers have already informally witnessed collective intelligence [35, 36] described in Section 3, often referred to as “emergent” [37] to describe situations where simple, seemingly blind individual behaviors aggregate to a complex and coherent behavior of the group. 2 shows that, when the individual bees’ behavior is modeled as $R_{\text{wvoter}}$, this “emergent” collective intelligence is MCL, a multi-armed bandit Reinforcement Learning algorithm. But the implications of 2 go even further: not only a swarm of bees is a single MCL agent, it is one that is quite efficient at what it does, because each bee is a parallel action sample of its policy. In other words, each bee $i$ can be seen as an action sample $a^{(i)}\sim\pi$ tested against a parallel copy $E^{(i)}$ of the environment $E$, as illustrated in Fig. 1. Fig. 2(a) shows that this converges much faster than if individual bees were themselves to learn via $\alpha$-MCL instead of just following $R_{\text{wvoter}}$. If the MCL algorithm were of any use outside of biology [^2] Most likely not: more advanced algorithms exist to solve multi-armed bandits (such as UCB [9])., say if a computer scientist wanted to train a robot via iterative $\alpha$-MCL, they would need a small $\alpha$ for the algorithm to converge to the optimal option. Or, if feasible, they could afford using a larger $\alpha$ by evaluating the policy in several parallel simulations (see LABEL:app:streaming_rl and LABEL:app:parallel_rl). Swarms of honey bees naturally learn via parallel RL in as many environments as there are scouts in the swarm. Finally, 1 implies that a similar discussion holds for the “imitation of success” revision protocol, which encompasses certain human behaviors, and even natural selection itself.
4.2 Swarm size
@@
In Section 4.1, we have discussed how being made of a number of individuals enables the “hive mind” to evaluate actions sampled from $\pi$ in a parallel fashion. Furthermore, our analysis assumed that this number was large (which enabled us to compute flows in Section 3). Yet, in practice there is a tradeoff here. Actual honey bee swarms are known to use only a small proportion of their bees as scouts, while the large majority of the swarm remains quiescent during nest-hunting [38, 39], which is surmised to be in the interest of energy-efficiency. In a typical swarm of 10,000 bees, roughly 200 to 500 scouts are active during nest-site selection. To analyze how the number of scouts $N$ affects the task-performance of the MCL macro-agent, we conducted simplified simulations where swarms of varying sizes had to choose the best amongst 10 options. Option qualities were spread between 0 and 1, while the noise in bees’ estimates was modeled with a uniform perturbation of amplitude $0.2$. Fig. 2(b) reports the percentage of populations that converged to the optimal choice for different values of $N$, across 1000 seeds (each individual seed converged to a homogeneous choice). In these simulations, $N=500$ is enough to closely follow the MRD and converge to the optimal option, whereas a number of scout bees that is too small often yields convergence to a sub-optimal nest site option. However, our choice of simulation parameters is largely arbitrary: Fig. 2(b) should be interpreted qualitatively rather than quantitatively.
4.3 Neighborhood size
In Section 3, we have assumed that the local neighborhood size $M$ was large. This assumption was useful to analytically derive the macro-agent RL update, because it meant that local neighborhoods (i.e., other scout bees that a scout may randomly walk into within its vicinity) could be considered well-mixed. Arguably, actual neighborhood sizes are not that large in the real world. Therefore, to complete our theoretical analysis, we performed $R_{\text{wvoter}}$ simulations with varying neighborhood sizes. Fig. 2(c) shows that even a small $M\geq 5$ yields a macro-agent algorithm that closely follows MCL. With $M=1$, each scout copies an arbitrary neighbor, which yields no macro-dynamic.
4.4 The collective power of promoting
@@
The $R_{\text{wvoter}}$ revision protocol adopted by honey bees has surprising advantages over imitation of success. In LABEL:sup:methods, we show that $R_{\text{wvoter}}$ (MCL) generally converges faster than $R_{\text{success}}$ (CL), except for small population sizes. This is in addition to the simplicity of $R_{\text{wvoter}}$, which only requires bees to blindly mimic their peers.
4.5 Could bees do more?
In Section 3, we showed that the bees employing $R_{\text{wvoter}}$ can be seen as a single online RL agent. However, a natural question arises: is $R_{\text{wvoter}}$ the optimal collective-decision making strategy? To answer this question, we investigate three simple variants of the imitation of success model and the weighted voter model (more details in LABEL:app:pop_details):
  • Deterministic Imitation of Success: Unlike the standard imitation of success model (which switches stochastically based on partner’s success), this variant deterministically adopts the partner’s type when it has more success.
  • Deterministic Imitation of Success with Weighted Voter Rule: This variant combines the weighted voter model’s success-weighted neighbor sampling with the deterministic imitation of success switch based on comparative success.
  • Stochastic Imitation of Success with Weighted Voter Rule: This variant differs from Deterministic Imitation of Success with Weighted Voter Rule in that it uses classic stochastic imitation of success instead of our deterministic version as the switching rule.
Fig. 2(d) shows that, at least in this simplified environment, all these variants converge to the optimal decision faster than $R_{\text{wvoter}}$ or MRD. This simple experiment suggests that there may be better collective decision-making strategies than $R_{\text{wvoter}}$. In particular, we see that if the bees could perform comparisons with their individual quality estimates before blindly imitating others from the neighborhood, the convergence speed could be faster than MRD. This insight raises an important question: why did bees evolve to use $R_{\text{wvoter}}$ [19, 23] over other collective decision-making strategies?
We speculate that this behaviour relates to cognitive constraints and the fundamental challenge of quality comparison. The weighted voter model requires remarkably minimal individual capabilities: bees need only estimate the quality of their own option and observe the types (not quality estimates) of neighbors, broadcasting their preference at a frequency proportional to quality. $R_{wvoter}$ requires no comparison operations between quality values and no direct communication of numerical quality estimates, capabilities that all faster-converging variants demand. In contrast, deterministic switching rules require bees to not only observe others’ quality estimates and compare them to their own, but also presume an objective or at least mutually calibrated scale for evaluating nest site quality. This is non-trivial: individual bees may experience different environmental conditions during their assessments, have different sensory acuity, or weight various quality features differently. $R_{wvoter}$ may elegantly sidestep this calibration problem by converting subjective quality estimates into broadcast frequencies — a transformation that preserves relative preferences while eliminating the need for interpersonal comparison. The weighted voter model may thus represent an evolutionary solution that achieves collective reinforcement learning with the absolute minimum individual cognitive machinery while avoiding the difficult problem of quality scale calibration, trading some convergence speed for dramatic simplification of individual processing and robustness to individual variation.
This hypothesis aligns with the broader pattern in social insects where complex collective behaviors emerge from remarkably simple individual rules, suggesting that evolutionary pressures favor cognitive parsimony at the individual level when collective intelligence can compensate at the group level, a strategy that may be driven in part by the substantial metabolic costs of neural tissue, which in mammals requires nearly an order of magnitude more energy per unit weight than other somatic tissues [40, 41]. We leave this exploration open for future work.
@@ 4.6 Broader significance
  • Biology. 2 shows that seemingly blind, imitation-based bee interactions aggregate to a form of collective intelligence, which is that of a “hive mind” learning through reinforcement. This is interesting from a neuroscience perspective, since both imitation and reinforcement have long been known to govern learning at the individual level [42, 4]. What our paper shows about honey bees is that their puzzling imitative behavior can in fact be understood as reinforcement from the swarm-agent perspective, where this behavior equates to massively parallel MCL. Why this imitative behavior has evolved in nature may thus appear clearer: swarms of honey bees are an example of reinforcement strategy having evolved at the group level. This analysis agrees with the commonly accepted theory of group/kin selection to explain the behavior of honey bees [43]. Building on the framework established in 1, parallel findings in other phylogenetically distant biological species underscore the generality of the proposed population–policy equivalences of this work. Notably, pheromone-guided C. elegans foraging exhibits the same mathematical structure as cross-learning [44], suggesting that stigmergic coordination may also be modeled as a form of collective reinforcement learning.
  • Economy and Social Dynamics. Evolutionary Game Theory goes beyond modeling biology: it extends to many types of population dynamics, most notably in economics [33, 29]. When thinking of $R_{\text{success}}$ as a model of certain decision-making strategies, 1 yields that imitating others’ success makes us part of a macroscopic RL process that hugely benefits the whole group, by mutualizing our speed of convergence to optimal strategies (see Section 4.1). This opens an avenue for modeling the group-level impacts of economic actors’ imitative behaviors as forms of collective reinforcement learning.
  • Algorithmic Swarm Intelligence. The field of Swarm Intelligence takes inspiration from complex emergent behaviors arising from a collective of natural entities following simple, local, and decentralized rules [35], to engineer algorithms that leverage collective principles like coordination, cooperation, and communication to tackle various problems (including CDM) across multiple domains such as swarm robotics [45, 46] and optimization [47]. Propositions 1 and 2 provide theoretical grounding for the phenomenon commonly referred to as “emergence” in this field [37]. This bridge also opens up the possibility of porting ideas from Reinforcement Learning to derive Swarm Intelligence local rules.
  • Algorithmic Reinforcement Learning. Similar to how we abstracted a population of $R_{wvoter}$ individuals as a single MCL agent, other revision models (majority rule [19] and cross-inhibition [48]) can be used to abstract dynamic groups of entities as simple, single RL agents. In massively multi-agent real-world settings (e.g., finance), this may greatly help model the non-stationarity of environment dynamics.
@@ 4.7 Conclusion
We have demonstrated that imitation-based collective behavior in large populations can be mathematically equivalent to reinforcement learning, providing a unifying framework that reinterprets swarm intelligence, social dynamics, and evolutionary processes as emergent forms of collective learning. This result provides a formal basis for understanding how simple, local imitation behaviors, often regarded as non-cognitive, can give rise to group-level intelligence and adaptive learning at scale. In particular, 2 demonstrates that the “hive mind” observed in swarms of bees can indeed be viewed as a single coherent agent learning via reinforcement, giving rise to a new bandit algorithm that we coin Maynard-Cross Learning.